# A Buddy for Betelgeuse: Binarity as the Origin of the Long Secondary Period in Orionis

### v1

@@ Abstract
We predict the existence of Ā OriĀ B, a low-mass companion orbiting Betelgeuse. This is motivated by the presence of a 2170-day Long Secondary Period (LSP) in Betelgeuseās lightcurve, a periodicity times longer than the starās 416 day fundamental radial pulsation mode. While binarity is currently the leading hypothesis for LSPs in general, the LSP and the radial velocity variation observed in Betelgeuse, taken together, necessitate a revision of the prevailing physical picture. The lightcurve-RV phase difference requires a companion to be behind Betelgeuse at the LSP luminosity minimum, 180 degrees out of phase with the system orientation associated with occultation. We demonstrate the consistency of this model with available observational constraints and identify tensions in all other proposed LSP hypotheses. Within this framework, we calculate a mass for Ā OriĀ B of and an orbital separation of , or times the radius of Betelgeuse. We then describe the features of the companion as constrained by the fundamental parameters of Betelgeuse and its orbital system, and discuss what would be required to confirm the companionās existence observationally.
@@ 1 Introduction
The idea that binarity could explain LSPs dates back to the late 1990s and early 2000s (Wood etĀ al., 1999a, 2004b) and has been reinforced since by the sequences formed by variable stars on period-luminosity diagrams (SoszyÅski, 2007) and in RV data (Hinkle etĀ al., 2002; Nicholls etĀ al., 2009). While such early versions of the binarity hypothesis were most concerned with whether a close companion could induce low-frequency modes on the primary star, the current leading theory is that the timescale of the LSP period is set by the orbital time of a low-mass companion, and the mechanism of dimming and brightening involves the formation and removal of dust along the line of sight in phase with the companionās orbit (SoszyÅski etĀ al., 2021).
@@ 2 LSPs in Cool Evolved Stars
As shown in Figure 1, LSPs in sequence D also follow a period-luminosity relationship that is distinct from the P-L relations discussed so far. Sequence D was discovered by Wood etĀ al. (1999a) using MACHO data. Binarity was proposed as a possible mechanism for the LSP by various authors (Hinkle etĀ al., 2002; Olivier & Wood, 2003; Wood etĀ al., 2004b). This was reaffirmed when Derekas etĀ al. (2006) showed that sequence D is an extension of sequence E (eclipsing binaries).
These sequences can also be extended further into the red supergiant regime. Kiss etĀ al. (2006) found that Galactic supergiants form sequences that correspond to the O1, FM and LSP periodicities and extend from sequences B, C, and D, respectively. We recalculated the absolute brightnesses of this RSG sample, using Gaia DR3 distances from the catalog of Bailer-Jones etĀ al. (2021) and plot these as blue filled circles in FigureĀ 1. These distances were sufficient in most cases, but a few very bright stars required other sources for parallax (van Leeuwen, 2007; Reid etĀ al., 2019). For Betelgeuse, we used the distance calculated by Joyce etĀ al. (2020), shown as red filled circles.
Compared to the original plot by Kiss etĀ al. (2006), the O1 and FM sequences reproduced in FigureĀ 1 appear to be considerably tighter and nicely follow sequences B and Cā, which correspond to the O1 and FM modes in RGB stars. The LSP periods show larger scatter and have larger uncertainties, but appear to form a coherent extension of sequence D. The periods identified in Betelgeuse as the O1, FM and LSP align with these sequences.
In the case of radial pressure modes, in which the pulsation period is determined by the stellar radius and the average sound speed, a P-L relation is expected, as the period scales with average density, thereby radius and mass, and thereby luminosity. In this case, the longest possible radial pressure mode is the FM. The apparent existence of a P-L relation in the LSP, on the other hand, is puzzling prima facie. Gravity modes can yield periods below the FM, but no such period-luminosity relation exists for g-modes.
@@
It is reasonable to ask why something without an intrinsic source of variability would nonetheless display a P-L relation. For this we can look outside the surface of the star, and make a dynamical argument. Generically, for a virial fluid element or body outside the stellar radius, a period is set by the Keplerian orbital timescale at that location. Although a generic orbit is not sensitive to the stellar luminosity, any orbital architecture knows about the stellar mass and radius, as where is some characteristic distance from the center of the star to the point external to the stellar surface from which the LSP is modulated.
@@ 3 Characterizing the Lightcurve-RV Phase Offset
An important diagnostic for the nature of pulsational versus orbital behavior is the presence or lack of radial velocity (RV) modulations and the phase relative to brightness fluctuations. Figure 2 shows the recent lightcurve (LC) and RV variations for Betelgeuse side by side. The photometric data include V-band observations collected by the AAVSO as well as the processed and rectified photometry of the star from the SMEI instrument (Jackson etĀ al., 2004; Hick etĀ al., 2007), originally published by Joyce etĀ al. (2020). RV observations were collected for more than a decade with the STELLA robotic telescope at Tenerife (Granzer etĀ al., 2022). From the latter, an LSP period of Ā d and a peak-to-peak amplitude of km/s were determined.
@@
where is the average brightness or velocity, is the common frequency between the two data sets, and and are the Fourier amplitude and phase components, respectively. For simplicity, we limit the fit to . However, we calculate the phase offset in Betelgeuse not just for the LSP but for the FM and O1 modes as well, based on the -band and SMEI urves from Joyce etĀ al. (2020), truncated to the interval of the RV observations.
Figure 3 shows the phase offset between Betelgeuseās lightcurve and RV curve in the context of similar stars measured with Gaia. We compare Betelgeuse to a limited sample of LPV-candidate Gaia stars from Gaia Collaboration etĀ al. (2023a) for which we could estimate the phase offsets. We also show the pulsational phase lags for classical Cepheids for comparison, including high-precision data for 87 stars from the work of Szabó etĀ al. (2007), and lower precision data from Gaia DR3 for about 800 stars for which phase data are included in the database directly (Ripepi etĀ al., 2023).
@@ 4.1 Hypothesis 1: LSP is actually the FM
Joyce etĀ al. (2020) provided evidence that the 2100-day periodicity cannot be the fundamental mode on the basis that a periodicity of this length is not activated by the -mechanism in either linear oscillation calculations performed with GYRE (Townsend & Teitler, 2013; Townsend etĀ al., 2018) or 1D hydrodynamic calculations performed with MESA (Paxton etĀ al., 2011, 2013, 2015, 2018, 2019; Jermyn etĀ al., 2023). Likewise, MacLeod etĀ al. (2023) demonstrate that since the Great Dimming, Betelgeuseās fundamental mode oscillation was perturbed into a combination of overtones. As in Joyce etĀ al. (2020), MacLeod etĀ al. (2023) identify the 400-d periodicity as the FM by performing adiabatic and non-adiabatic calculations with GYRE, exploring a larger range of azimuthal orders than in Joyce etĀ al. (2020) (see their Table 1). Similarly, Jadlovský etĀ al. (2024) examine the behavior of shockwaves in the aftermath of the dimming, coming to the same conclusion as MacLeod etĀ al. (2023) that the dominant pulsation mode had transitioned from the FM to the first overtone.
@@ 4.2 Hypothesis 2: Giant convective cells
In the extended convective envelopes of RGB stars, AGB stars, and RSGs, indeed . In such a star, large in the interior yields large-scale plumes spanning a sizeable fraction of the stellar radius, as recovered both by simulations (e.g. Freytag etĀ al., 2002; Dorch, 2004; Brun & Palacios, 2009; Plez & Chiavassa, 2013; Chiavassa etĀ al., 2009, 2010b, 2010a, 2011a, 2011b, 2017, 2024; Goldberg etĀ al., 2022) and inferred from observations (e.g. MontargĆØs etĀ al., 2017; Chiavassa etĀ al., 2018a, b; Kravchenko etĀ al., 2018, 2019, 2020, 2021; Zhang etĀ al., 2024). For RSGs like Betelgeuse, MLT predicts convective velocities of km/s for RSG stars (see discussions in Joyce etĀ al., 2020; Goldberg etĀ al., 2022), somewhat consistent with the required convective velocity to give rise to an overturn time of period in a star of radius
@@ 4.4.1 Differential Rotation
On the main sequence, the cores of high-mass stars rotate about 100x as fast as their surfaces (see, e.g., Heger etĀ al. 2000). However, the relationship between envelope and core rotation along the main sequence is not necessarily preserved once the star enters the core helium-burning phase. Mosser etĀ al. (2024) find that, for low-mass red giants, the evolution of the core rotation rate in core-helium-burning stars scales with the inverse square of the stellar radius, in good agreement with Tayar etĀ al. (2019). Although Betelgeuse is a massive star, and the core rotation rates of red supergiants have not been probed directly, asteroseismic relations for low-mass, core helium-burning red giants are informative for this case. The model described in Mosser etĀ al. (2024) is that of a rigid, dense helium core surrounded by a radiative layer with smooth rotation profile, all encased by a rigid convective envelope: a picture that describes Betelgeuse reasonably well. Putting the Joyce etĀ al. (2020) mass () and radius () for Betelgeuse into Mosser etĀ al. (2024)ās equation 3 yields a core rotation rate of approximately 400 days. This is much shorter than the LSP, so modulation by internal rotational is ruled out.
Though Lovekin etĀ al. (2009) note that extreme cases of differential rotation will cause slight shifts in acoustic mode frequencies, the extent of such a shift would be nowhere near the factor of decrease in frequency required to make the FM mimic the the LSP. Further, Jermyn etĀ al. (2020) find that there is very little radial differential rotation in stars with convective envelopes (in both short and long period eclipsing binaries) anyway.
@@ 4.5 Hypothesis 5: Magnetism
Some magnetic phenomena affecting high-mass stars are worth considering. The interplay between differential rotation and magnetic fields has been studied primarily in OB stars (Maeder & Meynet, 2011, 2012), which as Betelgeuseās spectral type when it was on the main sequence. (Keszthelyi etĀ al., 2019) find that fossil fields in massive stars can induce mass-loss quenching and magnetic braking; however, episodic mass loss is periodic and should happen at the frequency of the FM. The role of fossil fields would be to suppress the amount of mass dispersed per episode, not to alter the frequency of mass-loss episodes. Fossil fields therefore do not cause the LSP. Finally, while shorter-period, dynamo-generated magnetic fields can exist in the sub-surface convection zones in hot, massive stars (Cantiello & Braithwaite, 2011), Betelgeuse is not hot enough to have an iron opacity peak near the surface.
@@ 4.7 Hypothesis 7: -mechanism in the dust shell
where is the average density and the second proportionality assumes the mass scale is dominated by the stellar material and not the dust.111In order to achieve a modulated RV signal, near stellar-photosphere material would need to take part in the oscillation, so the the stellar mass is expected to set the mean density unless there is very appreciable mass within the dust shell. For an LSP:FM period ratio of , assuming the FM is an oscillation at the (optical or IR) stellar photosphere, this would mean that the āouterā radius mediating the RV signal would be times the stellar radius, or for a radius of (Joyce etĀ al., 2020). This is consistent with placement inside the region probed by radio observations indicating the presence of significant circumstellar dust from stellar radii (Altenhoff etĀ al., 1979; Lim etĀ al., 1998; Harper etĀ al., 2001; Smith etĀ al., 2009; OāGorman etĀ al., 2015).

### v2

# A Buddy for Betelgeuse: Binarity as the Origin of the Long Secondary Period in Ā Orionis

@@ Abstract
We predict the existence of Ā OriĀ B, a low-mass companion orbiting Betelgeuse. This is motivated by the presence of a 2170-day Long Secondary Period (LSP) in Betelgeuseās lightcurve, a periodicity times longer than the starās 416 day fundamental radial pulsation mode. While binarity is currently the leading hypothesis for LSPs in general, the LSP and the radial velocity variation observed in Betelgeuse, taken together, necessitate a revision of the prevailing physical picture. Specifically, the lightcurve-RV phase difference requires a companion to be behind Betelgeuse at the LSP luminosity minimum, 180 degrees out of phase with the system orientation associated with occultation. We demonstrate the consistency of this model with available observational constraints and identify tensions in all other proposed LSP hypotheses. Within this framework, we calculate a mass for Ā OriĀ B of and an orbital separation of , or times the radius of Betelgeuse. We then describe the features of the companion as constrained by the fundamental parameters of Betelgeuse and its orbital system, and discuss what would be required to confirm the companionās existence observationally.
@@ 1 Introduction
The idea that binarity could explain LSPs dates back to the late 1990s and early 2000s (Wood etĀ al., 1999a, 2004b) and has been reinforced since by the sequences formed by variable stars on period-luminosity diagrams (SoszyÅski, 2007) and in RV data (Hinkle etĀ al., 2002; Nicholls etĀ al., 2009). While such early versions of the binarity hypothesis were most concerned with whether a close companion could induce low-frequency modes on the primary star, the current leading theory is that the timescale of the LSP period is set by the orbital time of a low-mass companion, and the mechanism of dimming and brightening involves the formation and removal of dust along the line of sight in phase with the companionās orbit (SoszyÅski etĀ al., 2021). Additionally, it has been proposed before that Betelgeuse is a multiple system (Karovska etĀ al., 1986) from speckle-imaging measurements, inferred to be two high-eccentricity companions including one at a separation of 0.06ā with a period of 2.1 years, which has not been confirmed in follow-up work with higher-resolution instruments (e.g. Wilson etĀ al., 1992; Kervella etĀ al., 2009; MontargĆØs etĀ al., 2016).
@@ 2 LSPs in Cool Evolved Stars
LSPs in sequence D also follow a period-luminosity relationship that is distinct from the P-L relations discussed so far. Sequence D was discovered by Wood etĀ al. (1999a) using MACHO data. Binarity was proposed as a possible mechanism for the LSP by various authors (Hinkle etĀ al., 2002; Olivier & Wood, 2003; Wood etĀ al., 2004b). This was reaffirmed when Derekas etĀ al. (2006) showed that sequence D is an extension of sequence E (eclipsing binaries).
These sequences can also be extended further into the RSG regime (also shown in FigureĀ 1). Kiss etĀ al. (2006) found that Galactic supergiants form sequences that correspond to the O1, FM and LSP periodicities and extend from sequences B, C, and D, respectively. We recalculated the and absolute brightnesses of this RSG sample, using Gaia DR3 distances from the catalog of Bailer-Jones etĀ al. (2021) and plot these as blue filled circles in FigureĀ 1. These distances were sufficient in most cases, but a few very bright stars required other sources for parallax (van Leeuwen, 2007; Reid etĀ al., 2019). For Betelgeuse, we used the distance calculated by Joyce etĀ al. (2020), shown as red filled circles. Here we show Wesenheit indices as a proxy for luminosity based on the (J-K) colors, using the equation: , because that limits the effects of interstellar reddening and helps to separate the sequences better (Lebzelter etĀ al., 2019).
Compared to the original plot by Kiss etĀ al. (2006), the O1 and FM sequences reproduced in FigureĀ 1 appear to be considerably tighter. However, mode assignment based on PL or PW relations can be complicated as the constituent stars may belong to very different evolutionary stages, luminosity regimes, and variability classes that follow different relations. To illustrate this, we include three different PW relations with blue dashed lines, computed by SoszyÅski etĀ al. (2007), who identified further sequences and substructures in the red giant sample. The PW relation running through sequence C belongs to the Mira stars, whereas a1 and b1 belong to the lower-luminosity AGB and RGB stars, respectively. The FM supergiants lie close to the extrapolations of these relations. We discuss modal assignment in more detail and based on further arguments in Sect.Ā 4.1.
The RSG LSP periods in FigureĀ 1 show larger scatter and have larger uncertainties, but appear to form a coherent extension of sequence D. The periods identified in Betelgeuse as the O1, FM and LSP align with the RSG sequences. In the case of radial pressure modes, in which the pulsation period is determined by the stellar radius and the average sound speed, a P-L relation is expected, as the period scales with average density, thereby radius and mass, and thereby luminosity. In this case, the longest possible radial pressure mode is the FM. The apparent existence of a P-L relation in the LSP, on the other hand, is puzzling prima facie. Gravity modes can yield periods below the FM, but no such period-luminosity relation exists for g-modes.
@@
It is reasonable to ask why something without an intrinsic source of variability would nonetheless display a P-L relation. For this we can look outside the surface of the star, and make a dynamical argument. Generically, for a virial fluid element or body outside the stellar radius, a period is set by the Keplerian orbital timescale at that location. Although a generic orbit is not sensitive to the stellar luminosity, any orbital architecture knows about the stellar mass and radius, as , where is the distance from the center of the star to the external location setting the long secondary period timescale, i.e. the orbital separation.
@@ 3 Characterizing the Lightcurve-RV Phase Offset
An important diagnostic for the nature of pulsational versus orbital behavior is the presence or lack of radial velocity (RV) modulations and the phase relative to brightness fluctuations.
Figure 2 shows the recent lightcurve (LC) and RV variations for Betelgeuse side by side. The photometric data include V-band observations collected by the American Association of Variable Star Observers (AAVSO) as well as the processed and rectified photometry of the star from the SMEI instrument (Jackson etĀ al., 2004; Hick etĀ al., 2007), originally published by Joyce etĀ al. (2020). RV observations were collected for more than a decade with the STELLA robotic telescope at Tenerife (Granzer etĀ al., 2022). From the latter, an LSP period of Ā d and a peak-to-peak amplitude of km/s were determined.
@@
where is either the average brightness or the average (also called ) velocity, is the common frequency between the two data sets, and and are the Fourier amplitude and phase components, respectively. For simplicity, we limit the fit to . However, we calculate the phase offset in Betelgeuse not just for the LSP but for the FM and O1 modes as well, based on the -band and SMEI curves from Joyce et al. (2020), plus newer -band observations collected by the AAVSO, truncated to the interval of the RV observations. We list our results for Betelgeuse in Table 1. The light curve and pulsation periods of Betelgeuse are known to vary, and they went through considerable changes after the Great Dimming again, with O1 becoming the dominant pulsation mode instead of the FM (Dupree et al., 2022; Jadlovský et al., 2024). With the incorporation of new observations, the average periods of both modes shifted slightly relative to the values found by Joyce et al. (2020): the O1 became longer, whereas the FM became shorter, matching the period found by Kiss et al. (2006).
Figure 3 shows the phase offsets between Betelgeuseās lightcurve and RV curve in the context of similar stars measured with Gaia. We compare Betelgeuse to a limited sample of LPV-candidate Gaia stars from Gaia Collaboration etĀ al. (2023a) for which we could estimate the phase offsets. We also show the pulsational phase lags for classical Cepheids for comparison, including high-precision data for 87 stars from the work of Szabó etĀ al. (2007), and lower precision data from Gaia DR3 for about 800 stars for which phase data are included in the database directly (Ripepi etĀ al., 2023).
@@ 4.1 Hypothesis 1: LSP is actually the FM
Joyce etĀ al. (2020) provided evidence that the 2100-day periodicity cannot be the fundamental mode on the basis that a periodicity of this length is not activated by the -mechanism in either linear oscillation calculations performed with GYRE (Townsend & Teitler, 2013; Townsend etĀ al., 2018) or 1D hydrodynamic calculations performed with MESA (Paxton etĀ al., 2011, 2013, 2015, 2018, 2019; Jermyn etĀ al., 2023). Likewise, MacLeod etĀ al. (2023) demonstrate that since the Great Dimming, Betelgeuseās fundamental mode oscillation was perturbed into a combination of overtones. As in Joyce etĀ al. (2020), MacLeod etĀ al. (2023) identify the 400-d periodicity as the FM by performing adiabatic and non-adiabatic calculations with GYRE, exploring a larger range of spherical degrees than in Joyce etĀ al. (2020) (see their Table 1). Similarly, Jadlovský etĀ al. (2024) examine the behavior of shockwaves in the aftermath of the dimming, coming to the same conclusion as MacLeod etĀ al. (2023) that the dominant pulsation mode had transitioned from the FM to the first overtone.
@@ 4.2 Hypothesis 2: Giant convective cells
In the extended convective envelopes of RGB stars, AGB stars, and RSGs, indeed . In such a star, large in the interior yields large-scale plumes spanning a sizeable fraction of the stellar radius, as recovered both by simulations (e.g. Freytag etĀ al., 2002; Dorch, 2004; Brun & Palacios, 2009; Plez & Chiavassa, 2013; Chiavassa etĀ al., 2009, 2010b, 2010a, 2011a, 2011b, 2017, 2024; Goldberg etĀ al., 2022; Antoni & Quataert, 2022) and inferred from observations (e.g. MontargĆØs etĀ al., 2017; Chiavassa etĀ al., 2018a, b; Kravchenko etĀ al., 2018, 2019, 2020, 2021; Zhang etĀ al., 2024). For RSGs like Betelgeuse, MLT predicts convective velocities of km/s for RSG stars (see discussions in Joyce etĀ al., 2020; Goldberg etĀ al., 2022), somewhat consistent with the required convective velocity to give rise to an overturn time of period in a star of radius
@@
The MLT treatment which defines a single āconvective velocityā scale and thus a single āconvective overturnā timescale is, of course, a dramatic simplification of the true complexities of convection (see e.g. Joyce & Tayar 2023 §5, as well as the discussion in §5 of Goldberg etĀ al. 2022 in the case of RSGs, especially their Fig.Ā 19). Stars like Betelgeuse with large envelopes hosting vigorous convection are indeed more poorly approximated by MLT than solar-like stars by virtue of their significant deviation from solar conditions (see e.g. Joyce & Chaboyer 2018). Locally Super-Eddington luminosities due to the presence of strong opacity peaks lead to numerical as well as physical challenges for 1D models, which must be overcome through various āengineeringā techniques (see, e.g., §7.7 of Paxton etĀ al. 2013 and §7.2 of Jermyn etĀ al. 2023; and discussions and references therein). Moreover, the convective motions observed in 3D Radiation-Hydrodynamics simulations of near-surface convection in RSG envelopes other luminous massive stars do not even resemble motions typical of MLT-like convection out to the surface (Schultz etĀ al., 2020, 2022; Goldberg etĀ al., 2022); above some radius () the correlations between the radial fluid velocity and the density, entropy, and opacity invert ā cold, dense, high-opacity material which would sink in MLT-like convection is correlated with positive ā even at high optical depth .111 While there are significant caveats with MLT in the context of defining a single persistent convective overturn time, it is unlikely that the simplifying assumptions of MLT discussed above would affect results in this work in any other way. As noted by Joyce etĀ al. (2020), the primary impact of changing in models of 10-25 was to shift in temperature the location of the red giant branch, leading to a factor-of-ten increase in the uncertainty of the effective temperature constraint as compared to the error bars reported observationally by Levesque & Massey (2020). While uncertainty in the optimal value of the mixing length could certainly impact results through ās impact on the synthetic seismic frequency spectra against which Betelgeuseās pulsation spectrum was compared, this uncertainty was fully accounted for in Joyce etĀ al. (2020), whose parameters we use elsewhere in our analysis.
@@ 4.4.1 Differential Rotation
On the main sequence, the cores of high-mass stars have been predicted by some models to rotate about 100x as fast as their surfaces (see, e.g., Heger etĀ al. 2000, 2005), though more recent studies have shown that these models over-predicted core rotation by a factor of 10. In one such study, Pedersen (2022) found that measurements indicate that SPB stars with masses of 5-10 rotate close to rigidly. Conversely, Hermes etĀ al. (2017) do find evidence of a connection between high stellar birth mass and rapid core rotation via asteroseismology of white dwarfs, although again for masses below that of Betelgeuse. There are currently no measurements of core rotation in stars above , so extrapolation to Betelgeuse should be taken with caution.
Further, the relationship between envelope and core rotation along the main sequence is not necessarily preserved once the star enters the core helium-burning phase. Mosser etĀ al. (2024) find that, for low-mass red giants, the evolution of the core rotation rate in core-helium-burning stars scales with the inverse square of the stellar radius, in good agreement with Tayar etĀ al. (2019). Although Betelgeuse is a massive star, and the core rotation rates of red supergiants have not been probed directly, asteroseismic relations for low-mass, core helium-burning red giants are informative for this case. The model described in Mosser etĀ al. (2024) is that of a rigid, dense helium core surrounded by a radiative layer with smooth rotation profile, all encased by a rigid convective envelope: a picture that describes Betelgeuse reasonably well. Putting the Joyce etĀ al. (2020) lower mass bound () and radius () for Betelgeuse into Mosser etĀ al. (2024)ās equation 3 yields a core rotation rate of approximately 400 days. This is much shorter than the LSP, so modulation by internal rotational is ruled out.
Though Lovekin etĀ al. (2009) note that extreme cases of differential rotation will cause slight shifts in acoustic mode frequencies, the extent of such a shift would be nowhere near the factor of decrease in frequency required to make the FM mimic the LSP. Further, Jermyn etĀ al. (2020) find that there is very little radial differential rotation in stars with convective envelopes (in both short and long period eclipsing binaries) anyway.
@@ 4.5 Hypothesis 5: Magnetism
Some magnetic phenomena affecting high-mass stars are worth considering. The interplay between differential rotation and magnetic fields has been studied primarily in OB stars (Maeder & Meynet, 2011, 2012), which was Betelgeuseās spectral type when it was on the main sequence. (Keszthelyi etĀ al., 2019) find that fossil fields in massive stars can induce mass-loss quenching and magnetic braking; however, episodic mass loss is periodic and should happen at the frequency of the FM. The role of fossil fields would be to suppress the amount of mass dispersed per episode, not to alter the frequency of mass-loss episodes. Fossil fields therefore do not cause the LSP. Finally, while shorter-period, dynamo-generated magnetic fields can exist in the sub-surface convection zones in hot, massive stars (Cantiello & Braithwaite, 2011), Betelgeuse is not hot enough to have an iron opacity peak near the surface.
@@ 4.7 Hypothesis 7: -mechanism in the dust shell
where is the average density and the second proportionality assumes the mass scale is dominated by the stellar material and not the dust.222In order to achieve a modulated RV signal, near stellar-photosphere material would need to take part in the oscillation, so the stellar mass is expected to set the mean density unless there is very appreciable mass within the dust shell. For an LSP:FM period ratio of , assuming the FM is an oscillation at the (optical or IR) stellar photosphere, this would mean that the āouterā radius mediating the RV signal would be times the stellar radius, or for a radius of (Joyce etĀ al., 2020). This is consistent with placement inside the region probed by radio observations indicating the presence of significant circumstellar dust from stellar radii (Altenhoff etĀ al., 1979; Lim etĀ al., 1998; Harper etĀ al., 2001; Smith etĀ al., 2009; OāGorman etĀ al., 2015).